# CRAFT: Continuous Reasoning and Agentic Feedback Tuning for Multimodal Text-to-Image Generation

### v1

@@ 1 Introduction
Recent progress in large language models has demonstrated that explicit inference-time reasoning–including verification, reflection, and targeted correction–can substantially improve reliability without additional training. Rather than relying solely on larger models or additional data, these systems benefit from structured forms of thinking that make intermediate decisions explicit, checkable, and correctable. Related ideas have recently begun to influence multimodal generation, where reflection-based and agentic pipelines show that post-hoc critique can improve text-to-image (T2I) outputs without retraining [zhuo2025reflection, liu2025maestro].
At the same time, modern T2I systems remain predominantly single-shot and reactive. Despite impressive visual fidelity achieved by diffusion- and transformer-based models [rombach2022high, ramesh2022hierarchical, saharia2022photorealistic, podell2023sdxl, betker2023improving, yu2022scaling], they often struggle with long, compositional, and text-heavy prompts, where multiple constraints must be satisfied simultaneously. Generation is typically controlled through a single textual channel–the prompt–which acts as the primary mechanism for steering content, composition, style, and even safety in black-box deployments.
While inference-time refinement can mitigate some of these issues, existing approaches often rely on implicit, holistic critiques or unconstrained prompt rewrites [pryzant2023automatic, zhuo2025reflection, liu2025maestro], making their behavior difficult to interpret, control, or stop reliably. A key challenge is that image quality in open-domain T2I is inherently multi-criteria: a satisfactory image must satisfy compositional constraints, maintain object attributes and relations, adhere to style and layout, render text correctly, and avoid visual artifacts–all at once.
Recent work therefore emphasizes explicit, question-based evaluation, where prompts are decomposed into verifiable visual statements that better correlate with human judgments [hu2023tifa, lin2024evaluating]. Benchmarks such as Davidsonian Scene Graph (DSG) and GenEval further demonstrate that structured, checkable constraints provide a more reliable foundation for evaluating and improving T2I systems [cho2024davidsonian, ghosh2023geneval]. These insights suggest that open-domain T2I models could similarly benefit from converting implicit textual requirements into explicit, verifiable questions.
@@ 3 CRAFT T2I Generator
Visual-question generation: We construct a set of explicit visual questions $\mathcal{Q}$ that describe the intended scene and, when applicable, a dependency graph $\mathcal{G}$ so that prerequisite questions precede their dependents (e.g., “Is there a cat?” before “Is the cat red?”) [cho2024davidsonian]. The base question set covers objects, attributes, spatial relations, and high-level scene properties derived from the user prompt. In addition, we augment $\mathcal{Q}$ with two generic diagnostic queries: a text-visibility question “Is the text in the attached image fully clear and easily readable by a human?” and an artifact check “Does the image look natural and free from visual glitches, artifacts, or distortions?” Both are answered strictly with YES/NO to keep the signal discrete. The text-visibility question is included only when the prompt contains explicit textual content that is intended to appear in the image (e.g., spans in quotation marks or other text-to-render instructions). In such cases, the question acts as a global check on whether the rendered typography is legible, complementing the fine-grained DVQ questions about whether specific strings appear. The artifact check is always present and serves as a generic quality probe, capturing issues such as distorted text, warped geometry, or rendering glitches. Together, these additions extend the question set beyond purely semantic correctness to cover text legibility and low-level visual fidelity, which are then used as constraints in subsequent refinement steps.
Prompt rewrite: use an LLM to rewrite $\mathcal{P}_{0}$ following best practices for T2I prompting, reducing ambiguity while preserving the original intent [pryzant2023automatic]. At each iteration $t$, the system applies a controlled rewriting step to produce an improved prompt while preserving the user’s intent. The rewriting module is implemented as a constrained LLM-based editor that operates under a fixed set of rules. First, the module enforces a canonical structure for text-to-image prompting – explicitly ensuring that the rewritten prompt contains a well-formed subject description, contextual/environmental details, and a stylistic specification. If the original prompt does not specify a style, the model inserts a default realistic style. The editor also enhances semantic clarity by adding descriptive modifiers and resolving ambiguous phrasing, while avoiding stylistic drift beyond the bounds of the user’s initial intent.
@@
These rationales guide the next prompt edits [zhuo2025reflection] for better generation results.
@@ 5 Dataset and Models
We evaluated using DSG and DVQ metrics [cho2024davidsonian]. In addition to the main 2-iteration configuration, we tested multiple backbone models and multiple iteration budgets for both generation and editing (see Model configurations for baseline and CRAFT). For each configuration, we computed all metrics across several datasets to assess robustness and cross-model consistency. For text-to-image (T2I), we ran a 2-iteration setup with the Qwen-Image model [wu2025qwen]. For image editing, we used a 2-iteration setup with Gemini 2.5 Flash by Google. We further included two preference-based metrics inspired by the Maestro paper – auto side-by-side and auto side-by-side advantage [liu2025maestro]. As evaluation data, we used DSG-1K [cho2024davidsonian] and the Parti-Prompt (P2) suite [yu2022scaling]. Results are presented below (Table 1, Table 2).
@@ 6.1 Model Selection
  • FLUX-Schnell and FLUX-Dev: Open-source diffusion models with strong baseline quality [flux2025].
  • Qwen-Image: Multimodal autoregressive model with native text-to-image generation [wu2025qwen].
  • Z-Image-Turbo: Novelly published, open-source diffusion model with fast inference speed and one of the best baseline quality  [cai2025zimage].
@@ 6.3 System Prompts
  • DSG Question Generation: Constructs visual questions and dependency graphs from prompts, as it was implemented in [cho2024davidsonian].
@@ 8.1 Results on DSG1K Dataset
Table 1: Results on DSG1K [cho2024davidsonian] dataset.

### v2

@@ 1 Introduction
Recent progress in large language models has demonstrated that explicit inference-time reasoning–including verification, reflection, and targeted correction–can substantially improve reliability without additional training. Rather than relying solely on larger models or additional data, these systems benefit from structured forms of thinking that make intermediate decisions explicit, checkable, and correctable. Related ideas have recently begun to influence multimodal generation, where reflection-based and agentic pipelines show that post-hoc critique can improve text-to-image (T2I) outputs without retraining [19, 16].
At the same time, modern T2I systems remain predominantly single-shot and reactive. Despite impressive visual fidelity achieved by diffusion- and transformer-based models [14, 13, 15, 11, 1, 18], they often struggle with long, compositional, and text-heavy prompts, where multiple constraints must be satisfied simultaneously. Generation is typically controlled through a single textual channel–the prompt–which acts as the primary mechanism for steering content, composition, style, and even safety in black-box deployments.
While inference-time refinement can mitigate some of these issues, existing approaches often rely on implicit, holistic critiques or unconstrained prompt rewrites [12, 19, 16], making their behavior difficult to interpret, control, or stop reliably. A key challenge is that image quality in open-domain T2I is inherently multi-criteria: a satisfactory image must satisfy compositional constraints, maintain object attributes and relations, adhere to style and layout, render text correctly, and avoid visual artifacts–all at once.
Recent work therefore emphasizes explicit, question-based evaluation, where prompts are decomposed into verifiable visual statements that better correlate with human judgments [7, 8]. Benchmarks such as Davidsonian Scene Graph (DSG) and GenEval further demonstrate that structured, checkable constraints provide a more reliable foundation for evaluating and improving T2I systems [4, 5]. These insights suggest that open-domain T2I models could similarly benefit from converting implicit textual requirements into explicit, verifiable questions.
@@ 3 CRAFT T2I Generator
Visual-question generation: We construct a set of explicit visual questions $\mathcal{Q}$ that describe the intended scene and, when applicable, a dependency graph $\mathcal{G}$ so that prerequisite questions precede their dependents (e.g., “Is there a cat?” before “Is the cat red?”) [4]. The base question set covers objects, attributes, spatial relations, and high-level scene properties derived from the user prompt. In addition, we augment $\mathcal{Q}$ with two generic diagnostic queries: a text-visibility question “Is the text in the attached image fully clear and easily readable by a human?” and an artifact check “Does the image look natural and free from visual glitches, artifacts, or distortions?” Both are answered strictly with YES/NO to keep the signal discrete. The text-visibility question is included only when the prompt contains explicit textual content that is intended to appear in the image (e.g., spans in quotation marks or other text-to-render instructions). In such cases, the question acts as a global check on whether the rendered typography is legible, complementing the fine-grained DVQ questions about whether specific strings appear. The artifact check is always present and serves as a generic quality probe, capturing issues such as distorted text, warped geometry, or rendering glitches. Together, these additions extend the question set beyond purely semantic correctness to cover text legibility and low-level visual fidelity, which are then used as constraints in subsequent refinement steps.
Prompt rewrite: use an LLM to rewrite $\mathcal{P}_{0}$ following best practices for T2I prompting, reducing ambiguity while preserving the original intent [12]. At each iteration $t$, the system applies a controlled rewriting step to produce an improved prompt while preserving the user’s intent. The rewriting module is implemented as a constrained LLM-based editor that operates under a fixed set of rules. First, the module enforces a canonical structure for text-to-image prompting – explicitly ensuring that the rewritten prompt contains a well-formed subject description, contextual/environmental details, and a stylistic specification. If the original prompt does not specify a style, the model inserts a default realistic style. The editor also enhances semantic clarity by adding descriptive modifiers and resolving ambiguous phrasing, while avoiding stylistic drift beyond the bounds of the user’s initial intent.
@@
These rationales guide the next prompt edits [19] for better generation results.
@@ 5 Dataset and Models
We evaluated using DSG and DVQ metrics [4]. In addition to the main 2-iteration configuration, we tested multiple backbone models and multiple iteration budgets for both generation and editing (see Model configurations for baseline and CRAFT). For each configuration, we computed all metrics across several datasets to assess robustness and cross-model consistency. For text-to-image (T2I), we ran a 2-iteration setup with the Qwen-Image model [17]. For image editing, we used a 2-iteration setup with Gemini 2.5 Flash by Google. We further included two preference-based metrics inspired by the Maestro paper – auto side-by-side and auto side-by-side advantage [16]. As evaluation data, we used DSG-1K [4] and the Parti-Prompt (P2) suite [18]. We also report results on the P2-hard subset of Parti-Prompt, following the Maestro protocol. Results are presented below (Table 1, Table 3).
@@ 6.1 Model Selection
  • FLUX-Schnell and FLUX-Dev: Open-source diffusion models with strong baseline quality [2].
  • Qwen-Image: Multimodal autoregressive model with native text-to-image generation [17].
  • Z-Image-Turbo: Novelly published, open-source diffusion model with fast inference speed and one of the best baseline quality  [3].
@@ 6.3 System Prompts
  • DSG Question Generation: Constructs visual questions and dependency graphs from prompts, as it was implemented in [4].
@@ 6.4 Computational Cost
Latency.
A single CRAFT iteration (generation or editing) takes $\approx 30$ seconds on average in our setup; three iterations typically yield strong results. In contrast, a single Hunyuan Image 3.0 generation averages $\approx 60$ seconds while underperforming CRAFT-equipped mid-cost models on compositional quality.
@@ 8.1 Results on DSG1K Dataset
Table 1: Results on DSG1K [4] dataset.
@@
8.2 Comparison with Prompt Optimization Methods
Table 2 compares CRAFT with prompt-optimization baselines reported in Maestro [16]. We report only DSGScore and omit ranks. Note that our CRAFT results use a GPT-based VLM judge (gpt-5-nano-2025-08-07), whereas Maestro reports scores with Gemini 2.0 Flash, so absolute values are not directly comparable across papers.
Table 2: Comparison with prompt-optimization methods from Maestro. We report DSGScore only; higher is better.
Method P2-hard DSGScore $\uparrow$ DSG-1K DSGScore $\uparrow$
Original 0.826 0.772
Rewrite 0.855 0.815
Promptist [6] 0.873 0.849
LM-BBO [9] 0.859 0.806
OPT2I [10] 0.900 0.838
Maestro [16] 0.92 0.882
CRAFT (Ours) 0.90 0.91
CRAFT achieves the best DSGScore on DSG-1K and is competitive on P2-hard, where Maestro reports slightly higher scores. Overall, constraint-driven refinement provides stronger compositional gains than generic prompt-rewrite or agentic optimization methods.
@@ 9.2 Empirical Cost–Quality Analysis
Latency and iteration budget.
In our experiments, a single CRAFT iteration (generation or editing) takes on average $\approx 30$ seconds, and three iterations are typically sufficient to reach strong qualitative results. By comparison, a single Hunyuan Image 3.0 generation takes $\approx 60$ seconds on average while delivering lower compositional quality than CRAFT-equipped mid-cost backbones under our evaluation.